# A Comprehensive Assessment Benchmark for Rigorously Evaluating Deep Learning Image Classifiers

### v2

@@ Abstract
Reliable and robust evaluation methods are a necessary first step towards developing machine learning models that are themselves robust and reliable. Unfortunately, current evaluation protocols typically used to assess classifiers fail to comprehensively evaluate performance as they tend to rely on limited types of test data, and ignore others. For example, using the standard test data fails to evaluate the predictions made by the classifier to samples from classes it was not trained on. On the other hand, testing with data containing samples from unknown classes fails to evaluate how well the classifier can predict the labels for known classes. This article advocates bench-marking performance using a wide range of different types of data and using a single metric that can be applied to all such data types to produce a consistent evaluation of performance. Using such a benchmark it is found that current deep neural networks, including those trained with methods that are believed to produce state-of-the-art robustness, are extremely vulnerable to making mistakes on certain types of data. This means that such models will be unreliable in real-world scenarios where they may encounter data from many different domains, and that they are insecure as they can easily be fooled into making the wrong decisions. It is hoped that these results will motivate the wider adoption of more comprehensive testing methods that will, in turn, lead to the development of more robust machine learning methods in the future.
@@ 1.3 Terminology
OOD generalisation is concerned with classifying data where the distribution of the input (i.e., the characteristics of the images) has changed, but the distribution of output (i.e., the range of class labels associated with the images) is constant. In contrast OOD rejection is concerned with identifying data where the distribution of both the input and output has changed. The term “distribution” is therefore being used to refer to different distributions in each case. For OOD generalisation it is the distribution of the input, while for OOD rejection it is primarily the distribution of the output (the change in the input distribution is simply a side-effect of the change to new class labels). To avoid any possible confusion, OOD detection/rejection will hence-forth be referred to as “Unknown Class Rejection”. The data used to assess unknown class rejection will be subdivided into “novel” classes (images containing objects from classes unseen during training) and “unrecognisable” images (those containing no sematically meaningful information).
The term “distribution” when applied to the input is also poorly defined and, arguably, irrelevant when discussing IID/OOD generalisation. All data other than that included in the training set is OOD for the classifier (Guérin et al.,, 2022), so the distinction between IID and OOD data is qualitative rather than quantitative, and the boundary between in- and out-of distribution is necessarily arbitrary. Furthermore, the shift in distribution in image-space often has little correlation with the difficulty in correctly classifying an image. For example, two images of the same object taken from different viewpoints may have very little similarity in terms of corresponding pixels values, but the images may be easily identified as containing the same object. In contrast, it may be very difficult for the classifier to assign the same label to two images that are almost identical in image-space (e.g., a clean image and an adversarially perturbed version of that same image). Most importantly, how far an image is from the training distribution is irrelevant to the task of the classifier. Ideally, the classifier should be able to successfully recognise novel exemplars of the objects it was trained on however similar or dissimilar they are in appearance to the images used in training. In contrast an ideal classifier should reject images containing unknown objects even if such images are close to the training distribution. Hence, rather than using the terms IID and OOD, the remaining text will refer to such data as being from “known” classes. The broad class of known data will be sub-divided into different types of test data-set: “clean” data from the standard test set, “corrupt” for samples from the standard test set that have been manipulated to test for generalisation to changes in viewing conditions, and “adversarial” for samples that have been adversarially-perturbed.
@@ 1.5 Issues with Current Metrics
Performance on known classes (clean, corrupt, and adversarial samples) is typically evaluated by calculating the classification accuracy (or equivalently the percentage error), see Fig. 2 and Table 1, although many other metrics can also be used (Guo et al.,, 2023). However, if models are to be allowed to reject samples, to deal with those from unknown classes, then the performance on known samples should also be assessed taking into account that some samples may be rejected. On the other hand, models designed to perform unknown class rejection (Kirchheim et al.,, 2022; Chen et al.,, 2023; Cheng et al.,, 2023; Xu-Darme et al.,, 2023; Yang et al.,, 2022; Yang et al., 2023b, ; Lee et al.,, 2022), and adversarial defences that reject perturbed samples (Stutz et al.,, 2020; Ren et al.,, 2022; Lee et al.,, 2022; Chen et al.,, 2022), are typically evaluated purely in terms of how accurately samples are rejected and accepted. In other words, the task is considered to be that of binary classification (see Fig. 2 and Table 1). Hence, standard metrics for binary classification can be used, such as the False Positive Rate when the true positive rate is X% (FPR@X%), which assesses performance for a single rejection threshold, or Area Under the Receiver Operating Characteristic curve (AUROC) or Area Under the Precision-Recall curve (AUPR) both of which assess performance over varying rejection thresholds. An issue with FPR@X% is that separate thresholds are used for distinguishing each unknown data-set from the clean test data. Worse still, some rejection methods have hyper-parameters that are tuned separately on each unknown data-set (Zhang et al., 2023b, ). This is unrealistic, as in the real-world a single rejection criteria with fixed hyper-parameters would need to work to reject any unknown data, regardless of its origin (Shafaei et al.,, 2019; Feng et al.,, 2022). An issue with all standard unknown class rejection metrics is that the ability of the model to accurately classify the accepted samples is not evaluated. In some publications the clean accuracy of the model is evaluated, but this is done without any samples being rejected, and hence, does not reflect the performance that would be expected in the real-world where the same rejection criteria would be applied to all samples.
@@
Some metrics do attempt to evaluate classification performance in the presence of rejection. For example, in the context of evaluating accuracy on adversarial data, confidence-thresholded robust error (Stutz et al.,, 2020) measures the classification error for test examples that are not rejected, when the rejection threshold has been set so that 1% of correctly classified clean examples can be rejected. Using the metric robust risk with detection (Tramèr,, 2021) a classifier is accurate if it correctly classifies the clean samples, and either rejects or correctly classifies perturbed samples. However, with these metrics a high score can be obtained by a classifier that rejects a large proportion of the perturbed samples, even though these samples come from a known class and may be only perturbed imperceptibly (Chen et al.,, 2022). To address this issue, robustness with rejection (Chen et al.,, 2022) defines a variable perturbation threshold that can vary from zero to the adversarial attack budget. For test samples with a perturbation less than or equal to the threshold, both incorrect classification and rejection is counted as an error. For test samples with a perturbation greater than the threshold, incorrect classification is considered an error, but rejection is not counted as an error. When the threshold is greater or equal to the attack perturbation budget, robustness with rejection is equivalent to standard robust error. When the threshold is zero this metric is equivalent to robust risk with detection as proposed by Tramèr, (2021).
@@
2 Methods
2.1 Proposed Assessment Method
This article proposes to extend and modify the evaluation method proposed by Zhu et al., (2024), described in the preceding section, in the following ways:
@@ 2.1 Proposed Assessment Method
Extending the evaluation to consider other forms of generalisation. Zhu et al., (2024) evaluate performance using several data-sets including clean and corrupt data and data from unknown classes. In the current work, unknown class rejection is evaluated not only using images from novel classes, as in (Zhu et al.,, 2024), but also using unrecognisable images. Furthermore, performance is also evaluated using adversarial samples. In other words, here performance is evaluated using the five types of data illustrated in Fig. 1, whereas Zhu et al., (2024) use only three.
@@ 2.2 Choice of Test Data
The current work tested the performance of Deep Neural Networks (DNNs) trained on the following data-sets: CIFAR10, CIFAR100, TinyImageNet, and MNIST. In each case training was performed with the standard training sub-set of data. The robustness of each trained network was tested using the five categories of test data illustrated in Fig. 1. Within each category the following data was used.
@@
Popular common corruptions data-sets were used: CIFAR10-C, CIFAR100-C, TinyImageNet-C and MNIST-C (Hendrycks and Dietterich,, 2019; Mu and Gilmer,, 2019). The first three data-sets contain 18 different corruptions including different types of noise, blurring, synthetic weather conditions, and digital corruptions. MNIST-C contains 15 different corruptions including different types of noise, blurring, geometric transformations, and superimposed patterns. As typical in the literature, performance was evaluated using all the corruptions at all degrees of intensity.
@@
Samples were produced using AutoAttack (AA; Croce and Hein,, 2020), a state-of-the-art ensemble attack method that employs both gradient-based (white-box) and gradient-free (black-box) attacks. AA was implemented using the torchattacks PyTorch library (Kim,, 2021). Two sets of adversarial samples where created. Each set was created by perturbing the entire standard (clean) test-set, but with a different method of constraining the magnitude of the perturbation. Specifically, AA was used to apply both and -norm constrained attacks. The perturbation budget () used for each attack was the standard value used in the previous literature for each data-set. Specifically, was set to and 0.5 for and -norm attacks, respectively, against networks trained on the CIFAR10, CIFAR100, and TinyImageNet data-sets, and was set to 0.3 for -norm and to 2 for -norm attacks on MNIST trained networks.
@@
In each case, three data-sets containing objects of unknown classes were chosen. Specifically, the CIFAR10 and CIFAR100 trained networks were tested using novel classes from the test-sets of the Textures (Cimpoi et al.,, 2014) and SVHN (Netzer et al.,, 2011) data-sets and the CIFAR data-set that was not used for training (i.e., a CIFAR10 trained network was evaluated on its ability to reject samples from CIFAR100, and vice versa). For TinyImageNet trained networks novel class data came from the Textures test set (Cimpoi et al.,, 2014), the iNaturalist 2021 validation set (Van Horn et al.,, 2018), and the ImageNet-O data-set (Hendrycks et al., 2021c, ). The MNIST trained networks were tested using unknown, novel, classes from the test-sets of the Omniglot (Lake et al.,, 2015), FashionMNIST (Xiao et al.,, 2017) and KMNIST (Clanuwat et al.,, 2018) data-sets.
@@
Data to test unrecognisable image rejection was the same for all networks, and consisted of synthetic images randomly generated using four different methods. First, images containing random blobs, as used in (Hendrycks et al., 2019a, ). Secondly, images in which each pixel intensity value was independently and randomly selected from a uniform distribution. Thirdly, the images of the standard (clean) test set after a random permutation of all pixels. Fourthly, the images of the clean test set after randomising the phase, in the Fourier domain, of each image. Each of these four data-sets contained the same number of samples as the standard test-set associated with the training data.
@@
In principal, the amount of test data that could be used is unlimited (as there are countless images containing objects that any one classifier has not been trained to recognise, and due to the stochastic nature of the methods used to create corrupt, adversarial, and unrecognisable images). Hence, the above selection of test data necessarily represents a small, arbitrary, sub-set of all the possible data that could be used. In making this selection the aim was to define a reasonably wide variety of data while ensuring that evaluation could be performed in a reasonable time using reasonable resources. The results in Section 3 suggest that the chosen data is sufficiently challenging that modern DNNs are far from ceiling performance. However, as classification methods improve in the future, the data used for evaluation could be extended to allow a more thorough and challenging assessment in two ways: adding more data in each category, or by adding categories.
@@
Expanding the range of data in each of the existing categories could be done by: using new sets of clean data (such as CIFAR10.1 (Recht et al.,, 2018), or ImageNet-A (Hendrycks et al., 2021b, )); using different corruptions (e.g., different forms of noise or image transformation), higher magnitude corruptions (e.g., a wider range of scales such as used in the MNIST Large Scale data set (Jansson and Lindeberg,, 2020)), the different types of corruption used in the OOD-CV-v2 benchmark (Zhao et al.,, 2024), or objects shown against different backgrounds (as is used in the NICO challenge (Zhang et al., 2023b, )); using any of numerous alternative algorithms for generating adversarial images (Liang et al.,, 2022; Ren et al.,, 2020; Xu et al.,, 2020), or different methods of constraining the perturbations, or higher perturbation budgets; using the additional novel class data-sets proposed in the OpenOOD benchmark Yang et al., (2022); using fooling images (Kumano et al.,, 2022; Nguyen et al.,, 2015) that are unrecognisable images which have been optimised to be difficult to reject. Adversarial attack-type techniques could also be used not only to perturb clean test samples, as is currently the case, but samples from other data types to make these types of data more challenging to correctly process. New categories of test data that could be used to extend and increase the challenge of the proposed benchmark include: known classes presented in the form of drawings, art, or cartoons (such as in ImageNet-R (Hendrycks and Dietterich,, 2019) or PACS (Li et al.,, 2017)); or images where texture has been replaced by that from another image (Geirhos et al.,, 2019).

### v3

@@ Abstract
Reliable and robust evaluation methods are a necessary first step towards developing machine learning models that are themselves robust and reliable. Unfortunately, current evaluation protocols typically used to assess classifiers fail to comprehensively evaluate performance as they tend to rely on limited types of test data, and ignore others. For example, using the standard test data fails to evaluate the predictions made by the classifier to samples from classes it was not trained on. On the other hand, testing with data containing samples from unknown classes fails to evaluate how well the classifier can predict the labels for known classes. This article advocates benchmarking performance using a wide range of different types of data and using a single metric that can be applied to all such data types to produce a consistent evaluation of performance. Using the proposed benchmark it is found that current deep neural networks, including those trained with methods that are believed to produce state-of-the-art robustness, are vulnerable to making mistakes on certain types of data. This means that such models will be unreliable in real-world scenarios where they may encounter data from many different domains, and that they are insecure as they can be easily fooled into making the wrong decisions. It is hoped that these results will motivate the wider adoption of more comprehensive testing methods that will, in turn, lead to the development of more robust machine learning methods in the future.
@@ 1.3 Terminology
OOD generalisation is concerned with classifying data where the distribution of the input (i.e., the characteristics of the images) has changed, but the distribution of output (i.e., the range of class labels associated with the images) is constant. In contrast OOD rejection is concerned with identifying data where the distribution of both the input and output has changed. The term “distribution” is therefore being used to refer to different distributions in each case. For OOD generalisation it is the distribution of the input, while for OOD rejection it is primarily the distribution of the output (the change in the input distribution is simply a side-effect of the change to new class labels). To avoid any possible confusion, OOD detection/rejection will hence-forth be referred to as “Unknown Class Rejection”. The data used to assess unknown class rejection will be subdivided (as in Fig. 1) into “novel” classes (images containing objects from classes unseen during training) and “unrecognisable” images (those containing no sematically meaningful information).
The term “distribution” when applied to the input is also poorly defined and, arguably, irrelevant when discussing IID/OOD generalisation. All data other than that included in the training set is OOD for the classifier (Guérin et al.,, 2022), so the distinction between IID and OOD data is qualitative rather than quantitative, and the boundary between in- and out-of distribution is necessarily arbitrary. Furthermore, the shift in distribution in image-space often has little correlation with the difficulty in correctly classifying an image. For example, two images of the same object taken from different viewpoints may have very little similarity in terms of corresponding pixel values, but the images may be easily identified as containing the same object. In contrast, it may be very difficult for the classifier to assign the same label to two images that are almost identical in image-space (e.g., a clean image and an adversarially perturbed version of that same image). Most importantly, how far an image is from the training distribution is irrelevant to the task of the classifier. Ideally, the classifier should be able to successfully recognise novel exemplars of the objects it was trained on regardless of how similar or dissimilar they are in appearance to the images used in training. In contrast an ideal classifier should reject images containing unknown objects even if such images are close to the training distribution. Hence, rather than referring to samples as IID and OOD, the remaining text will refer to such data as being from “known” classes. The broad class of known data will be sub-divided (as in Fig. 1) into different types of test data-set: “clean” data from the standard test set, “corrupt” for samples from the standard test set that have been manipulated to test for generalisation to changes in viewing conditions, and “adversarial” for samples that have been adversarially-perturbed.
@@ 1.5 Issues with Current Metrics
Performance on known classes (clean, corrupt, and adversarial samples) is typically evaluated by calculating the classification accuracy (or equivalently the percentage error), see Fig. 2 and Table 1, although many other metrics can also be used (Guo et al.,, 2023). However, if models are to be allowed to reject samples, to deal with those from unknown classes, then the performance on known samples should also be assessed taking into account that some samples may be rejected. On the other hand, models designed to perform unknown class rejection (Kirchheim et al.,, 2022; Chen et al.,, 2023; Cheng et al.,, 2023; Xu-Darme et al.,, 2023; Yang et al.,, 2022; Yang et al., 2023b, ; Lee et al.,, 2022), and adversarial defences that reject perturbed samples (Stutz et al.,, 2020; Ren et al.,, 2022; Lee et al.,, 2022; Chen et al.,, 2022), are typically evaluated purely in terms of how accurately samples are rejected and accepted. In other words, the task is considered to be that of binary classification (see Fig. 2 and Table 1). Hence, standard metrics for binary classification can be used, such as the False Positive Rate when the true positive rate is X% (FPR@X%), which assesses performance for a single rejection threshold, or Area Under the Receiver Operating Characteristic curve (AUROC) or Area Under the Precision-Recall curve (AUPR) both of which assess performance over varying rejection thresholds. An issue with FPR@X% is that separate thresholds are used for distinguishing each unknown data-set from the clean test data. Worse still, some rejection methods have hyper-parameters that are tuned separately on each unknown data-set (Zhang et al., 2023b, ). This is unrealistic, as in the real-world a single rejection criteria with fixed hyper-parameters would need to work to reject any unknown data, regardless of its origin (Shafaei et al.,, 2019; Feng et al.,, 2022). An issue with all standard unknown class rejection metrics is that the ability of the model to accurately classify the accepted samples is not evaluated. In some publications the clean accuracy of the model is evaluated, but this is done without any samples being rejected, and hence, does not reflect the performance that would be expected in the real-world where the same rejection criteria would be applied to all samples. As a result of this issue, a model that could accurately reject unknown classes would score highly with the standard metrics, even if that classifier assigned a random label to all accepted samples making it virtually useless in any practical application.
@@
Some metrics do attempt to evaluate classification performance in the presence of rejection. For example, in the context of evaluating accuracy on adversarial data, confidence-thresholded robust error (Stutz et al.,, 2020) measures the classification error for test examples that are not rejected, when the rejection threshold has been set so that 1% of correctly classified clean examples can be rejected. Using the metric robust risk with detection (Tramèr,, 2021) a classifier is accurate if it correctly classifies the clean samples, and either rejects or correctly classifies perturbed samples. However, with these metrics a high score can be obtained by a classifier that rejects a large proportion of the perturbed samples, even though these samples come from a known class and may be only perturbed imperceptibly (Chen et al.,, 2022). To address this issue, robustness with rejection (Chen et al.,, 2022) defines a variable perturbation threshold that can vary from zero to the adversarial attack budget. For test samples with a perturbation less than or equal to the threshold, both incorrect classification and rejection is counted as an error. For test samples with a perturbation greater than the threshold, incorrect classification is considered an error, but rejection is not counted as an error. When the threshold is greater or equal to the attack perturbation budget, robustness with rejection is equivalent to standard robust error. When the threshold is zero this metric is equivalent to robust risk with detection as proposed by Tramèr, (2021), discussed earlier.
@@
2 Methods: The Proposed Comprehensive Assessment Benchmark
2.1 Assessment Method
The proposed benchmark uses a metric to evaluate performance that is an extended and modified version of the evaluation method proposed by Zhu et al., (2024), described in the preceding section. The proposed modifications are as follows:
@@ 2.1 Assessment Method
Extending the evaluation to consider other forms of generalisation. Zhu et al., (2024) evaluate performance using three types of test data: clean, corrupt, and novel. In the current work, unknown class rejection is evaluated not only using images from novel classes, as in (Zhu et al.,, 2024), but also using unrecognisable images. Furthermore, performance is also evaluated using adversarial samples. In other words, here performance is evaluated using the five types of data illustrated in Fig. 1, whereas Zhu et al., (2024) use only three. The results reported in Section 3 highlight the necessity of expanding the range of test data types used. The poorest performance of the tested networks is almost always for either adversarial samples or unrecognisable images, i.e., for the data-types that Zhu et al., (2024) do not use. Hence, the benchmark proposed by Zhu et al., (2024) is more likely to fail to identify robustness issues than the benchmark proposed here.
@@ 2.2 Choice of Test Data
| Training | Test Data | ||||||||||
| Data-set | Clean | Corrupt | Adversarial | Novel | Unrecognisable | ||||||
| MNIST | MNIST test-set | MNIST-C | AA | AA | Omniglot | FMNIST | KMNIST | Phase | Scramble | Blobs | Uniform |
| FMNIST | FMNIST test-set | FMNIST-C | AA | AA | Textures | MNIST | CIFAR10 | ||||
| SVHN | SVHN test-set | MNIST–C | AA | AA | Textures | Icons50 | CIFAR10 | ||||
| CIFAR10 | CIFAR10 test-set | CIFAR10-C | AA | AA | Textures | SVHN | CIFAR100 | ||||
| CIFAR100 | CIFAR100 test-set | CIFAR100-C | AA | AA | Textures | SVHN | CIFAR10 | ||||
| TinyIN | TinyIN val.-set | TinyIN-C | AA | AA | Textures | iNaturalist | ImageNetO | ||||
| ImageNet1k | ImageNet1k val.-set | ImageNet-C | AA | AA | Textures | iNaturalist | ImageNetO | ||||
The benchmark is defined to assess the performance of classifiers trained on seven popular image classification data-sets: MNIST (LeCun et al.,, 1998), FashionMNIST (FMNIST; Xiao et al.,, 2017), SVHN (Netzer et al.,, 2011), CIFAR10 (Krizhevsky,, 2009), CIFAR100 (Krizhevsky,, 2009), TinyImageNet (TinyIN), and ImageNet1k (a. k. a. ILSVRC2012; Russakovsky et al.,, 2015). Performance is evaluated using test-data from each of the five categories illustrated in Fig. 1. The specific data-sets used in each category is summarised in Table 2 and described in more detail below.
@@
Common corruptions data-sets: MNIST-C, FashionMNIST-C (FMNIST-C), CIFAR10-C, CIFAR100-C, TinyImageNet-C and ImageNet-C (Hendrycks and Dietterich,, 2019; Mu and Gilmer,, 2019; Weiss and Tonella,, 2022). These data-sets modify the standard test-set using multiple corruption methods including different types of noise, blurring, geometric transformations, superimposed patterns, synthetic weather conditions, and digital corruptions. As typical in the literature, performance is evaluated using all the corruptions at all degrees of intensity. When the test-set resolution differs from that of the training set, the test images is resized (using bi-linear interpolation) to match the image resolution used for training. Similarly, the number of colour channels in the test images is modified, if necessary, to match the training data (e.g., grayscale images are converted to colour by making the three colour channels all equal to the original image intensity values).
@@
Samples are produced using AutoAttack (AA; Croce and Hein,, 2020), a state-of-the-art ensemble attack method that employs both gradient-based (white-box) and gradient-free (black-box) attacks. AA is implemented using the torchattacks PyTorch library (Kim,, 2021). Two sets of adversarial samples are used. Each set is created by perturbing the entire standard (clean) test-setbbbDo to the extremely high computational demands of performing AA on large ImageNet1k trained networks, only 1000 adversarial samples were used for these networks., but with a different method of constraining the magnitude of the perturbation. Specifically, AA is used to apply both and -norm constrained attacks. The perturbation budget () used for each attack is the standard value used in the previous literature for each data-set, and is defined in Table 2.
@@
Three data-sets containing objects of unknown classes are used for each training data-set. Specifically, MNIST trained networks are tested using unknown, novel, classes from the test-sets of the Omniglot (Lake et al.,, 2015), FashionMNIST (Xiao et al.,, 2017) and KMNIST (Clanuwat et al.,, 2018) data-sets. FashionMNIST trained networks are evaluated with the test-sets of the Textures (Cimpoi et al.,, 2014), MNIST and CIFAR10 data-sets. When trained with SVHN, unknown class rejection is evaluated with the Textures, Icons-50 (Hendrycks and Dietterich,, 2019), and CIFAR10 test data-sets. The CIFAR10 and CIFAR100 trained networks are tested using novel classes from the test-sets of the Textures and SVHN data-sets and the CIFAR data-set that was not used for training (i.e., a CIFAR10 trained network is evaluated on its ability to reject samples from CIFAR100, and vice versa). For TinyImageNet and ImageNet1k trained networks novel class data comes from the Textures test set, the iNaturalist 2021 validation set (Van Horn et al.,, 2018), and the ImageNet-O data-set (Hendrycks et al., 2021c, ). When the test-set resolution differs from that of the training set, the test images is resized (using bi-linear interpolation) to match the image resolution used for training. Similarly, the number of colour channels in the test images is modified, if necessary, to match the training data (e.g., colour images are converted to grayscale).
@@
Data to test unrecognisable image rejection is the same for all training data-sets, and consisted of synthetic images randomly generated using four different methods. First, the images of the standard (clean) test set after randomising the phase, in the Fourier domain, of each image (“Phase”). Second, the images of the clean test set after a random permutation of all pixels (“Scramble”). Third, as used in Hendrycks et al., 2019a , images containing random blobs (“Blobs”). Fourth, images in which each pixel intensity value is independently and randomly selected from a uniform distribution (“Uniform”). Each of these four data-sets contain the same number of samples as the standard test-set associated with the training data.
@@
In principal, the amount of test data that could be used is unlimited (as there are countless images containing objects that any one classifier has not been trained to recognise, and due to the stochastic nature of the methods used to create corrupt, adversarial, and unrecognisable images). Hence, the above selection of test data necessarily represents a small, arbitrary, sub-set of all the possible data that could be used. In making this selection the aim was to define a reasonably wide variety of data while ensuring that evaluation could be performed in a reasonable time using reasonable resources. The results in Section 3 suggest that the chosen data is sufficiently challenging that modern Deep Neural Networks (DNNs) are far from ceiling performance. However, as classification methods improve in the future, the data used for evaluation could be extended to allow a more thorough and challenging assessment in two ways: adding more data in each category, or by adding categories.
@@
Expanding the range of data in each of the existing categories could be done by: using new sets of clean data, such as CIFAR10.1 (Recht et al.,, 2018), or ImageNet-A (Hendrycks et al., 2021b, ); using different corruptions (e.g., different forms of noise or image transformation), higher magnitude corruptions (e.g., a wider range of scales such as used in the MNIST Large Scale data set (Jansson and Lindeberg,, 2020)), the different types of corruption used in the OOD-CV-v2 benchmark (Zhao et al.,, 2024), or objects shown against different backgrounds (as is used in the NICO challenge (Zhang et al., 2023b, )); using any of numerous alternative algorithms for generating adversarial images (Liang et al.,, 2022; Ren et al.,, 2020; Xu et al.,, 2020), or different methods of constraining the perturbations, or higher perturbation budgets; using the additional novel class data-sets proposed in the OpenOOD benchmark (Yang et al.,, 2022); using fooling images (Kumano et al.,, 2022; Nguyen et al.,, 2015) that are unrecognisable images which have been optimised to be difficult to reject. Adversarial attack-type techniques could also be used not only to perturb clean test samples, as is currently the case, but samples from other data types to make these types of data more challenging to correctly process (Li et al.,, 2024). New categories of test data that could be used to extend and increase the challenge of the proposed benchmark include: known classes presented in the form of drawings, art, or cartoons, such as in ImageNet-R (Hendrycks and Dietterich,, 2019) or PACS (Li et al.,, 2017); or images where texture has been replaced by that from another image (Geirhos et al.,, 2019).